Calculate the national income and personal disposable income from the following information


1
Expert's answer
2021-07-05T17:31:03-0400

GNP = Y = Rs. 5000.00, Personal Disposable Income DI = 4100.00, Consumption C = 3800.00, X-M = 50.00, Govt. Budget Deficit BD = 200.00.

DI = Y - T, BD = G - T, Y = C + I + G + X - M, so: Government Expenditure G = BD + T = BD + (Y - DI) = 200 + 5000 - 4100 = 1100.

As Y = C + G + I + X - M, then Gross Investment I = Y - C - G - (X - M) = 5000 - 3800 - 1100 - 50 = 50.